okay this is a word problem that i dont know how to do:
Brad is 3 years more than half of Brandon's age. If their combined age is 93 years, how old is each man?

Answer by stanbon(75887)

Brad is 3 years more than half of Brandon's age. If their combined age is 93 years, how old is each man?
-----
Equations:
Let Brad's'age be "x":
Let Brandon's age be "y":
---
x = (y/2)+3
x + y = 93
----
Substitute for "x" and solve for "y":
(y/2) + 3 + y = 93
(3/2)y = 90
y = (2/3)90
y = 60 (Brandon's age now)
---
solve for "x":
x = (y/2)+3
x = 33 (Brad's age now)
=============
Cheers,
Stan H.
==============